# What can we learn from quantum convolutional neural networks?

### v2

@@ Abstract
We can learn from analyzing quantum convolutional neural networks (QCNNs) that: 1) working with quantum data can be perceived as embedding physical system parameters through a hidden feature map; 2) their high performance for quantum phase recognition can be attributed to generation of a very suitable basis set during the ground state embedding, where quantum criticality of spin models leads to basis functions with rapidly changing features; 3) pooling layers of QCNNs are responsible for picking those basis functions that can contribute to forming a high-performing decision boundary, and the learning process corresponds to adapting the measurement such that few-qubit operators are mapped to full-register observables; 4) generalization of QCNN models strongly depends on the embedding type, and that rotation-based feature maps with the Fourier basis require careful feature engineering; 5) accuracy and generalization of QCNNs with readout based on a limited number of shots favor the ground state embeddings and associated physics-informed models. We demonstrate these points in simulation, where our results shed light on classification for physical processes, relevant for applications in sensing. Finally, we show that QCNNs with properly chosen ground state embeddings can be used for fluid dynamics problems, expressing shock wave solutions with good generalization and proven trainability.
@@ I Introduction
In this work, we aim to demystify the inner workings of quantum convolutional neural networks, specifically highlighting why they are so successful in quantum phase recognition. We show that supplied quantum states (features) can be understood in terms of hidden feature maps — quantum processes that prepare states depending on classical parameters $\bm{x}$ (being a feature vector or scalar; see Fig. 1). During the mapping process, we observe that the ground state preparation supplies a very beneficial basis set, which allows the building of a nonlinear quantum model for the decision boundary — the ultimate goal of classification — that is sharp and generalizes on few data points and with smaller number of measurement shots. This is compared to the rotation-based Fourier embedding, which requires feature engineering for sufficient generalization. We connect the developed QCNN description with error correction and multi-scale entanglement renormalization ansatz (MERA) explanations from Ref. [75], and show that single-qubit observables can be used to “pick up” the most suitable basis functions, while leading to sampling advantage. Motivated by classification, we apply the described QCNN workflow with ground state embedding to solve regression problems motivated by fluid dynamics. This can offer QML tools to deal with critical phenomena with improved generalization.
@@ II.1 Quantum neural networks as Fourier-type models
Given this layout of quantum neural networks, the unsettling question arises: how does the Fourier model description fit the quantum data story and QCNN-based models that are too distinct from what we just described?
@@ II.2 Quantum convolutional neural networks: prior art
where the open boundary conditions are considered. This specific Hamiltonian is chosen as an example of non-trivial spin order in 1D, being uniquely related to the measurement-based quantum computing [89]. In Eq. (7) the first term corresponds to the spin-1/2 cluster Hamiltonian with three-body terms, the second term represents the transverse field, and the third term contains Ising interaction terms. The point $h_{\mathrm{x}}=J_{\mathrm{xx}}=0$ corresponds to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetric Hamiltonian that hosts in its ground state a symmetry protected topological (SPT) phase. For $J_{\mathrm{xx}}=0$ we can study the transition from SPT order at $J>h_{\mathrm{x}}$ to staggered ferromagnetic order at $J<h_{\mathrm{x}}$, with $h_{\mathrm{x}}/J=1$ corresponding to the critical point. In the SPT phase and for the open boundary the ground state is four-fold degenerate, and is characterized by the string order parameter and nonzero topological entanglement entropy [90]. Examples of QCNN-based classification show that by measuring the string order $\hat{\mathcal{O}}=\hat{Z}_{1}\hat{X}_{2}\hat{X}_{3}...\hat{X}_{N-2}\hat{X}_{N-1}\hat{Z}_{N}$ one can distinguish the SPT phase from other phases (open boundary). It is also known that string order implies the utility of states for performing measurement-based quantum computing [89] for which the ground state subspace (cluster states) represent the resource.
@@ III Model
Here, we take the QCNN structure for quantum phase recognition and connect it with the standard QML workflow. For this we note that quantum states $|\psi_{\alpha}\rangle$ are not abstract quantum “data points”, but in fact are the result of quantum processes that correspond to thermalization or ground state (GS) preparation. Specifically, they depend on properties of the underlying Hamiltonian, and its parameters representing features $x$. Namely, for the cluster-type Hamiltonian used in QCNNs so far these are couplings $J$, $J_{\mathrm{xx}}$ and magnetic field $h_{\mathrm{x}}$. These system parameters represent a feature vector $x=\{J,h_{\mathrm{x}},J_{\mathrm{xx}}\}$ (and each concrete realization is denoted by index $\alpha$). Even though we may not have access to them, they define the underlying behavior of the system we attempt to classify. The process of embedding these features as a part of the state preparation we call a hidden feature map (Fig. 1). The circuits for preparing QCNN input as ground states of models $|\psi_{0}(x)\rangle$ may differ, and can comes both from experiments (e.g. sensors) or specifically designed GS preparation schedule. In the following we assume that there exists a unitary $\hat{\mathcal{U}}_{\varphi}(x)$ or a completely positive trace-preserving map $\mathcal{E}_{\varphi}(x)$ for the ground state preparation of studied Hamiltonians, and summarize different options in the Appendix A.
@@ III.1 Cluster ground state embedding
Let us now consider an example that can shed light on the internal structure of quantum convolutional neural networks. For this we take a spin-1/2 cluster Hamiltonian with periodic boundary. In this case the ground state in the SPT phase is unique and non-degenerate. From a technical perspective this enables working at smaller system size without being harmed by finite size effects, and avoid thermal ensemble state preparation (as pure ground state suffices in this case). Specifically, we consider the Hamiltonian of the form
@@
We assume that $x$ changes from 0 to 1 via control of $h_{x}$, while $h_{z}$ is adjusted to keep $\varepsilon$ constant and small. Next, we prepare the ground state (GS) of ^$\hat{\mathcal{H}}(x)$ assuming one of the preparation strategies (Appendix A). The corresponding circuit, which we denote $\hat{\mathcal{U}}_{\varphi}(x)$, represents our feature map, acting such that $\hat{\mathcal{U}}_{\varphi}(x)|{\o}\rangle\eqqcolon|\psi_{0}(x)\rangle$ being the $x$-dependent ground state. Tuning $x$ from zero to one, we go from the SPT phase with the cluster GS to a trivial product state $|\psi_{0}(1)\rangle=|+\rangle^{\otimes N}$, through the critical point $x_{\mathrm{cr}}=1/2$ where the GS develops nontrivial correlations. We highlight that one of the possible preparation circuits can be based on the Hamiltonian variational ansatz (HVA) [91]. Due to the favorable dynamical Lie algebra scaling for this model [92] and large gradients for HVA in general [93], the corresponding feature map is efficient, meaning that the depth of the preparation circuit scales at most quadratically in the system size. However, for the purpose of numerical tests it is also instructive to use exact diagonalization, as it allows for clean studies of ground state embedding and its properties irrespective of variational preparation.
@@ III.2 Analysis of the ground state embedding from the QNN perspective
Once we have established the system Hamiltonian ^$\hat{\mathcal{H}}(x)$, the mapping procedure $\hat{\mathcal{U}}_{\varphi}(x)$, and the measurement (cost) operator $\hat{\mathcal{O}}$ [Eq. (11)], we proceed to classification from the point of view of quantum model building. For this, we study the basis set of our feature-dependent ground states $|\psi_{0}(x)\rangle$. Namely, we assume the embedded states to be decomposed in the computational basis $|\psi_{0}(x)\rangle\coloneqq\sum_{j}\phi_{j}(x)|j\rangle$, presented as computational $2^{N}$ basis states $\{|j\rangle\}$ with corresponding binary/integer representation, weighted by the $x$-dependent basis functions $\{\phi_{j}(x)\}_{j=0}^{2^{N}-1}$. It is instructive to visualize this basis in some form. We choose to project feature states onto the computational basis states (much like in case of QCBMs [45, 48]), reading out probabilities $|\phi_{j}(x)|^{2}=|\langle j|\psi_{0}(x)\rangle|^{2}$, which can be seen as diagonal matrix elements of the corresponding density operator $\hat{\rho}_{0}(x)=|\psi_{0}(x)\rangle\langle\psi_{0}(x)|$. We choose a minimal example with $N=3$, shown in Fig. 2(a), and plot the corresponding squared basis functions in Fig. 2(b). We observe that all basis functions undergo a drastic change exactly at the critical point $x_{\mathrm{cr}}$, introducing an inductive bias for building models with an inherent criticality. We observe different behavior is associated to matrix elements that involve ferromagnetic bitstrings $j=000$ and $j=111$, one-hot bistrings ($j=100,010,001$), and two-hot states $j=110,101,011$. Admittedly, these basis functions are shown in the computational basis, while the string order lies in the X Pauli plane. We next proceed to show how the required basis functions are “picked up” by the string order observable $\hat{\mathcal{O}}=(-1)^{3}\hat{X}_{1}\hat{X}_{2}\hat{X}_{3}$. This specific string contains only anti-diagonal matrix elements (of $-1$ entries), such that the opposite (i.e. bitwise conjugated $j\leftrightarrow\bar{j}$) pairs of states are connected. In Fig. 2(c) we plot the product of basis functions $\phi_{j}^{*}(x)\cdot\phi_{\bar{j}}(x)$ that are collected when estimating the expectation. One can see that all contributions represent sigmoid-like functions, with slightly increasing or decreasing fronts. Importantly, there are multiple degeneracies, such that one can pick multiple contributions when building the model (thus, success does not depend on specific projections and unique elements). Finally, we see that the sum of all antidiagonal products, being the expectation $\langle\psi_{0}(x)|\hat{\mathcal{O}}|\psi_{0}(x)\rangle$, represents a sharp decision boundary even at small system size. We proceed to see how this basis analysis plays out in the full QCNN workflow.
@@ III.3 Analysis of QCNN basis transformation from the QNN perspective
Finally, let us offer another understanding of the fixed QCNN workflow. The action of 9-3-1 QCNN leads to the model being $\langle\psi(x)|\hat{\mathcal{U}}_{\mathrm{QCNN}}^{\dagger}\hat{X}_{5}\hat{\mathcal{U}}_{\mathrm{QCNN}}|\psi(x)\rangle$, where $\hat{\mathcal{U}}_{\mathrm{QCNN}}$ collect all layers (notice that no mid-circuit measurements are required, and we can simply trace out anything but the middle qubit). We can then see the action of QCNN as measuring an effective “dressed” operator $\hat{\mathcal{U}}_{\mathrm{QCNN}}^{\dagger}\hat{X}_{5}\hat{\mathcal{U}}_{\mathrm{QCNN}}\cong\hat{X}_{1}\hat{X}_{2}...\hat{X}_{9}$, which is equal to the string order parameter up to the global phase. This simply means that pooling makes sure we pick up the relevant basis states from our embedding, while building the model on the small subset of qubits. This is very important for trainability, as we reduce the shot noise (smaller number of measurements is required for models based on single-qubit sampling), and still enjoy the access to the full basis. We believe that the very same strategy can be applied for building other QML models, and QCNN-like measurement adaptation can improve solvers beyond classification.
@@ IV.3 Generalization
Figure 6: (a) Decision boundaries for the two types of embeddings (ground state and Fourier rotation-based), for the finite number of shots. Shaded area shows the corresponding variance, where large variance leads to less accurate results. (b) Generalization measure as a function of the number of measurement shots.
@@
These results further demonstrate the benefit of the GS feature map. The sharp transition at criticality ensures that errors arising from access to a finite number of shots have a minimal impact on classification. We expect the separation in generalization between the two feature maps to be even more apparent if we also train the circuit with a finite number of shots for each expectation value estimation. This is likely to transform the smooth decision boundary into a noisy transition line, and the lower sharpness of the Fourier QCNN would be further exposed in this setting. Even more importantly, considering cases where data is limited (e.g. when only few runs of quantum hardware can be analyzed due to time constraints), the generalization on few training examples and read out with few samples can largely yield the classification for actual devices.
@@ IV.4 Solving regression tasks with QCNNs
We perform regression selecting $M=21$ equally-spaced training points over $[0,1]$, for both the ground state embedding based on cluster Hamiltonian (same as in the previous section) and the Fourier feature map based on rotations $\hat{R}^{i}_{\mathrm{Y}}(2\pi x)$ for each qubit $i$. Training is performed via Adam optimizer. The results are as shown in Fig. 7(a). We observe that the original QCNN with GS embeddings is able to represent the transition with high degree of generalization, and the Fourier model experiences oscillations while roughly representing the trend. We assign an accuracy metric for each model, corresponding to the percentage of test points for which the square error difference between the trained function and target function is with an error bound $\epsilon$. The test points are chosen as 100 uniformly spaced points and error bound as $\epsilon=0.05$. The accuracy of the QCNN model with cluster Hamiltonian embedding is then equal to 91%, mostly due to slightly different scaling of tails, and the Fourier-based QCNN has only 62% test accuracy due to poor generalization. We conclude that ground state based embedding with the known region of criticality and critical exponents can be translated to problem in other domains of physics, offering distinct model-building capabilities with good generalization.
@@
Overall, from the use of QCNNs for regression we can see that it is important for the basis functions from the choice of encoding to match the problem. This is particularly important for embedding choices based on ground states, where all the basis functions have a strong particular behaviour at the critical point. Additionally, one can make the GS feature map become variational as well, such that the critical region and sharpness are altered as needed. Once a suitable embedding choice is set, the QCNN can be used to represent the target function efficiently and learnt with proven trainability [80] — an important advantage over fully-connected quantum neural networks.

### v3

@@ Abstract
Quantum machine learning (QML) shows promise for analyzing quantum data. A notable example is the use of quantum convolutional neural networks (QCNNs), implemented as specific types of quantum circuits, to recognize phases of matter. In this approach, ground states of many-body Hamiltonians are prepared to form a quantum dataset and classified in a supervised manner using only a few labeled examples. However, this type of dataset and model differs fundamentally from typical QML paradigms based on feature maps and parameterized circuits. In this study, we demonstrate how models utilizing quantum data can be interpreted through hidden feature maps, where physical features are implicitly embedded via ground-state feature maps. By analyzing selected examples previously explored with QCNNs, we show that high performance in quantum phase recognition comes from generating a highly effective basis set with sharp features at critical points. The learning process adapts the measurement to create sharp decision boundaries. Our analysis highlights improved generalization when working with quantum data, particularly in the limited-shots regime. Furthermore, translating these insights into the domain of quantum scientific machine learning, we demonstrate that ground-state feature maps can be applied to fluid dynamics problems, expressing shock wave solutions with good generalization and proven trainability.
@@ I Introduction
In this work, we aim to demystify the inner workings of quantum convolutional neural networks, specifically highlighting why they are so successful in quantum phase recognition. We show that supplied quantum states (features) can be understood in terms of hidden feature maps — quantum processes that prepare states depending on classical parameters $\bm{x}$ (being a feature vector or scalar; see Fig. 1). During the mapping process, we observe that the ground state preparation supplies a very beneficial basis set, which allows the building of a nonlinear quantum model for the decision boundary — the ultimate goal of classification — that is sharp and generalizes on few data points and with smaller number of measurement shots. This is compared to the rotation-based Fourier embedding, which requires feature engineering for sufficient generalization. We note that the utility of ground state embeddings is not limited to QCNNs, but could be an avenue to improved performance for general quantum models. We connect the developed QCNN description with error correction and multi-scale entanglement renormalization ansatz (MERA) explanations from Ref. [75], and show that single-qubit observables can be used to “pick up” the most suitable basis functions, while leading to sampling advantage. Motivated by classification, we apply the described QCNN workflow with ground state embedding to solve regression problems motivated by fluid dynamics. This can offer QML tools to deal with critical phenomena with improved generalization.
@@ II.1 Quantum neural networks as Fourier-type models
Given this layout of quantum neural networks, an interesting question arises: how does the Fourier model description fit the quantum data story and QCNN-based models that are too distinct from what we just described?
@@ II.2 Quantum convolutional neural networks: prior art
where open boundary conditions are considered. This specific Hamiltonian was chosen as an example of non-trivial spin order in 1D, being uniquely related to measurement-based quantum computing [89]. In Eq. (7) the first term corresponds to the spin-1/2 cluster Hamiltonian with three-body terms, the second term represents the transverse field, and the third term contains Ising interaction terms. The point $h_{\mathrm{x}}=J_{\mathrm{xx}}=0$ corresponds to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetric Hamiltonian that hosts in its ground state a symmetry protected topological (SPT) phase. For $J_{\mathrm{xx}}=0$ one can study the transition from SPT order at $J>h_{\mathrm{x}}$ to staggered ferromagnetic order at $J<h_{\mathrm{x}}$, with $h_{\mathrm{x}}/J=1$ corresponding to the critical point. In the SPT phase and for the open boundary the ground state is four-fold degenerate, and is characterized by the string order parameter and nonzero topological entanglement entropy [90]. Examples of QCNN-based classification show that by measuring the string order $\hat{\mathcal{O}}=\hat{Z}_{1}\hat{X}_{2}\hat{X}_{3}...\hat{X}_{N-2}\hat{X}_{N-1}\hat{Z}_{N}$ one can distinguish the SPT phase from other phases (open boundary). It is also known that string order implies the utility of states for performing measurement-based quantum computing [89] for which the ground state subspace (cluster states) represent the resource. This model is interesting to study due to the several transitions that can be identified. However, from the QML perspective the introduction of critical lines makes the analysis of basis functions opaque, so we simplify the model in our following study.
@@ III Model
Here, we take the QCNN structure for quantum phase recognition and connect it with the standard QML workflow. For this we note that quantum states $|\psi_{\alpha}\rangle$ are not abstract quantum “data points”, but in fact are the result of quantum processes that correspond to thermalization or ground state (GS) preparation. Specifically, they depend on properties of the underlying Hamiltonian, and its parameters representing features $x$. Namely, for the cluster-type Hamiltonian used in QCNNs so far these are couplings $J$, $J_{\mathrm{xx}}$ and magnetic field $h_{\mathrm{x}}$. These system parameters represent a feature vector $x=\{J,h_{\mathrm{x}},J_{\mathrm{xx}}\}$ (and each concrete realization is denoted by index $\alpha$). Even though we may not have access to them, they define the underlying behavior of the system we attempt to classify. The process of embedding these features as a part of the state preparation we call a hidden feature map (Fig. 1). The circuits for preparing QCNN input as ground states of models $|\psi_{0}(x)\rangle$ may differ, and can come both from experiments (e.g. sensors) or specifically designed GS preparation schedule. In the following we assume that there exists a unitary $\hat{\mathcal{U}}_{\varphi}(x)$ or a completely positive trace-preserving map $\mathcal{E}_{\varphi}(x)$ for the ground state preparation of studied Hamiltonians, and summarize different options in the Appendix A.
@@ III.1 Cluster ground state embedding
Let us now consider an example that can shed light on the internal structure of quantum models working with quantum data. For this we take a spin-1/2 cluster Hamiltonian with periodic boundary. For understanding essential parts of the workflow, we prefer to set $J_{\mathrm{xx}}=0$ such that only one transition point and one variable function are analyzed. In this case the ground state in the SPT phase is unique and non-degenerate. From a technical perspective this enables working at smaller system size without being harmed by finite size effects, and avoid thermal ensemble state preparation (as pure ground state suffices in this case). Specifically, we consider the Hamiltonian of the form
@@
We assume that $x$ changes from 0 to 1 via control of $h_{x}$, while $h_{z}$ is adjusted to keep $\varepsilon$ constant and small (ensuring a finite but small symmetry breaking term). Next, we prepare the ground state (GS) of ^$\hat{\mathcal{H}}(x)$ assuming one of the preparation strategies (Appendix A). The corresponding circuit, which we denote $\hat{\mathcal{U}}_{\varphi}(x)$, represents our feature map, acting such that $\hat{\mathcal{U}}_{\varphi}(x)|{\o}\rangle\eqqcolon|\psi_{0}(x)\rangle$ being the $x$-dependent ground state. Tuning $x$ from zero to one, we go from the SPT phase with the cluster GS to a trivial product state $|\psi_{0}(1)\rangle=|+\rangle^{\otimes N}$, through the critical point $x_{\mathrm{cr}}=1/2$ where the GS develops nontrivial correlations. We highlight that one of the possible preparation circuits can be based on the Hamiltonian variational ansatz (HVA) [91]. Due to the favorable dynamical Lie algebra scaling for this model [92] and large gradients for HVA in general [93], the corresponding feature map is efficient, meaning that the depth of the preparation circuit scales at most quadratically in the system size. However, for the purpose of numerical tests it is also instructive to use exact diagonalization, as it allows for clean studies of ground state embedding and its properties irrespective of variational preparation.
@@ III.2 Analysis of the ground state embedding from the QNN perspective
Once we have established the system Hamiltonian ^$\hat{\mathcal{H}}(x)$, the mapping procedure $\hat{\mathcal{U}}_{\varphi}(x)$, and the measurement (cost) operator $\hat{\mathcal{O}}$ [Eq. (11)], we proceed to classification from the point of view of quantum model building. For this, we study the basis set of our feature-dependent ground states $|\psi_{0}(x)\rangle$. Namely, we assume the embedded states to be decomposed in the computational basis $|\psi_{0}(x)\rangle\coloneqq\sum_{j}\phi_{j}(x)|j\rangle$, represented as computational $2^{N}$ basis states $\{|j\rangle\}$ with corresponding binary/integer representation, weighted by the $x$-dependent basis functions $\{\phi_{j}(x)\}_{j=0}^{2^{N}-1}$. It is instructive to visualize this basis in some form. We choose to project feature states onto the computational basis states (much like in case of QCBMs [45, 48]), reading out probabilities $|\phi_{j}(x)|^{2}=|\langle j|\psi_{0}(x)\rangle|^{2}$, which can be seen as diagonal matrix elements of the corresponding density operator $\hat{\rho}_{0}(x)=|\psi_{0}(x)\rangle\langle\psi_{0}(x)|$. We choose a minimal example with $N=3$, shown in Fig. 2(a), and plot the corresponding squared basis functions in Fig. 2(b). We observe that all basis functions undergo a drastic change exactly at the critical point $x_{\mathrm{cr}}$, introducing an inductive bias for building models with an inherent criticality. We observe different behavior is associated to matrix elements that involve ferromagnetic bitstrings $j=000$ and $j=111$, one-hot bistrings ($j=100,010,001$), and two-hot states $j=110,101,011$. Admittedly, these basis functions are shown in the computational basis, while the string order lies in the X Pauli plane. We next proceed to show how the required basis functions are “picked up” by the string order observable $\hat{\mathcal{O}}=(-1)^{3}\hat{X}_{1}\hat{X}_{2}\hat{X}_{3}$. This specific string contains only anti-diagonal matrix elements (of $-1$ entries), such that the opposite (i.e. bitwise conjugated $j\leftrightarrow\bar{j}$) pairs of states are connected. In Fig. 2(c) we plot the product of basis functions $\phi_{j}^{*}(x)\cdot\phi_{\bar{j}}(x)$ that are collected when estimating the expectation. One can see that all contributions represent sigmoid-like functions, with slightly increasing or decreasing fronts. Importantly, there are multiple degeneracies, such that one can pick multiple contributions when building the model (thus, success does not depend on specific projections and unique elements). Finally, we see that the sum of all antidiagonal products, being the expectation $\langle\psi_{0}(x)|\hat{\mathcal{O}}|\psi_{0}(x)\rangle$, represents a sharp decision boundary even at small system size. We proceed to see how this basis analysis plays out in the full QCNN workflow.
@@ III.3 Analysis of QCNN basis transformation from the QNN perspective
Finally, let us offer another understanding of the fixed QCNN workflow. The action of 9-3-1 QCNN leads to the model being $\langle\psi(x)|\hat{\mathcal{U}}_{\mathrm{QCNN}}^{\dagger}\hat{X}_{5}\hat{\mathcal{U}}_{\mathrm{QCNN}}|\psi(x)\rangle$, where $\hat{\mathcal{U}}_{\mathrm{QCNN}}$ collect all layers (notice that no mid-circuit measurements are required, and we can simply trace out anything but the middle qubit). We can then see the action of QCNN as measuring an effective “dressed” operator $\hat{\mathcal{U}}_{\mathrm{QCNN}}^{\dagger}\hat{X}_{5}\hat{\mathcal{U}}_{\mathrm{QCNN}}\cong\hat{X}_{1}\hat{X}_{2}...\hat{X}_{9}$, which is equal to the string order parameter up to the global phase. This simply means that pooling makes sure we pick up the relevant basis states from our embedding, while building the model on the small subset of qubits. This is very important for trainability, as we reduce the shot noise (smaller number of measurements is required for models based on single-qubit sampling), and still enjoy the access to the full basis. We believe that the very same strategy can be extended beyond the QCNN and applied for building other QML models, and QCNN-like measurement adaptation can improve solvers beyond classification.
@@ IV.1 QCNN trained on cluster ground state embedding
It is important to note that even with an arbitrarily chosen variational ansatz, we can still get excellent results for classification using the GS feature map, as we demonstrate in Appendix C. This reinforces that it is the specific choice of feature map that provides the right basis for finding a sharp decision boundary.
@@ IV.3 Generalization
Figure 6: (a) Decision boundaries for the two types of embeddings (ground state and Fourier rotation-based), for a finite number of shots. Shaded area shows the corresponding variance, where large variance leads to less accurate results. (b) Generalization measure as a function of the number of measurement shots.
@@
These results further demonstrate the benefit of the GS feature map. The sharp transition at criticality ensures that errors arising from access to a finite number of shots have a minimal impact on classification. In Appendix D, we also show that the GS embedding maintains this advantage in learning a sharp decision boundary when considering a model both trained and tested with shot noise. The smooth decision boundary is transformed into a noisy transition line, and the lower sharpness of the Fourier QCNN is further exposed in this setting. Even more importantly, considering cases where data is limited (e.g. when only few runs of quantum hardware can be analyzed due to time constraints), the generalization on few training examples and read out with few samples can largely yield the classification for actual devices.
@@ IV.4 Solving regression tasks with QCNNs
We perform regression selecting $M=21$ equally-spaced training points over $[0,1]$, for both the ground state embedding based on cluster Hamiltonian (same as in the previous section) and the Fourier feature map based on rotations $\hat{R}^{i}_{\mathrm{Y}}(2\pi x)$ for each qubit $i$. Training is performed via Adam optimizer, and we return to the state vector simulation to compare the performance of the feature maps in the absence of shot noise. The results are as shown in Fig. 7(a). We observe that the original QCNN with GS embeddings is able to represent the transition with high degree of generalization, and the Fourier model experiences oscillations while roughly representing the trend. We assign an accuracy metric for each model, corresponding to the percentage of test points for which the square error difference between the trained function and target function is with an error bound $\epsilon$. The test points are chosen as 100 uniformly spaced points and error bound as $\epsilon=0.05$. The accuracy of the QCNN model with cluster Hamiltonian embedding is then equal to 91%, mostly due to slightly different scaling of tails, and the Fourier-based QCNN has only 62% test accuracy due to poor generalization. We conclude that ground state based embedding with the known region of criticality and critical exponents can be translated to problem in other domains of physics, offering distinct model-building capabilities with good generalization.
@@
Overall, from the use of QCNNs for regression we can see that it is important for the basis functions from the choice of encoding to match the problem. This is particularly important for embedding choices based on ground states, where all the basis functions have a strong particular behaviour at the critical point. This application of the QCNN with GS embedding to regression tasks is a form of transfer learning; the specific basis functions induced by the feature map allows the quantum model to perform particularly well for fitting a particular class of functions. Additionally, one can make the GS feature map become variational as well, such that the critical region and sharpness are altered as needed. Once a suitable embedding choice is set, the QCNN can be used to represent the target function efficiently and learnt with proven trainability [80] — an important advantage over fully-connected quantum neural networks.